# Geometric structure of shallow neural networks and constructive ${\mathcal{L}}^{2}$ cost minimization

### v2

@@ Abstract.
In this paper, we approach the problem of cost (loss) minimization in underparametrized shallow neural networks through the explicit construction of upper bounds, without any use of gradient descent. A key focus is on elucidating the geometric structure of approximate and precise minimizers. We consider shallow neural networks with one hidden layer, a ReLU activation function, an ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size $N>QM$ that can be arbitrarily large. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal to noise ratio of training inputs. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$. The proof of the upper bound yields a constructively trained network; we show that it metrizes a particular $Q$-dimensional subspace in the input space ${\mathbb{R}}^{M}$. We comment on the characterization of the global minimum of the cost function in the given context.
@@ 1. Introduction
Applications of shallow and deep neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite of those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation, [2, 6]. The common approach to the minimization of the cost (loss) function is based on the gradient descent flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios; but nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) remain quite mysterious.
In this paper, we address the cost minimization problem via explicit construction of upper bounds on the global minimum of the cost function, without invoking gradient descent at all. The inspiration for our approach is drawn from methods in mathematical physics aimed at determining the ground state energy for complex manybody quantum systems (such as large molecules), see for instance [9]. Our main goal here is to obtain a rigorous mathematical understanding of the geometric structure of (approximate) cost minimizers; we do not focus on the development of efficient algorithms, or exploration of application cases in the present work. To make our exposition broadly accessible, we will mostly employ standard mathematical terminology, while minimally referring to specialized nomenclatures specific to the neural networks literature.
The current paper is the first in a series of works in which we investigate the geometric structure of cost minimizers in neural networks (that is, weights and biases that minimize the cost). Here, we analyze underparametrized shallow neural networks in the context of supervised learning, with one hidden layer, a ReLU activation function, an ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size $N>QM$ that can be arbitrarily large - therefore, we are considering the underparametrized regime. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal to noise ratio of training inputs. First, we derive an upper bound on the minimum of the cost function with an explicit construction of suitable weights and biases. As a by-product of our proof, we obtain a constructively trained shallow network with explicit parameters. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$.
@@
The above assumptions allow for a succinct exposition of core geometric aspects of the problem; an analysis of more general situations is left for future work. For some thematically related background, see for instance [1, 7, 6, 8, 10, 11] and the references therein. We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in underparametrized Deep Learning (that is, multilayer) networks in a separate work, [4], which will invoke results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [3].
In this introductory section, we summarize the main results of this paper. We consider a shallow network for which $y_{j}\in{\mathbb{R}}^{Q}$ denotes the $j$-th output vector, and define the output matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q}\,,$ (1.1)
which we assume to be invertible. We assume the input space to be given by ${\mathbb{R}}^{M}$ with $M\geq Q$, and for $j=1,\dots,Q$, we let
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,.$ (1.2)
denote the matrix of training input vectors $x_{0,j,1}$, $i=1,\dots,N_{j}$, which belong to the output $y_{j}$. Letting $N:=\sum_{j=1}^{Q}N_{j}$, the full matrix of training inputs is given by
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (1.3)
We assume that $Q\leq M\leq MQ<N$ where $N$ can be arbitrarily large. We define the average of all training input vectors belonging to the output $y_{j}$,
for $j=1,\dots,Q$, and
$\Delta x_{0,j,i}:=x_{0,j,i}-\overline{x_{0,j}}\,.$ (1.5)
Moreover, we let
$\delta:=\sup_{i,j}|\Delta x_{0,j,i}|\,.$ (1.6)
We define
$\Delta X_{0,j}:=[\Delta x_{0,j,1}\cdots\Delta x_{0,j,i}\cdots\Delta x_{0,j,N_{j}}]\in{\mathbb{R}}^{M\times N_{j}}\,,$ (1.7)
and
$\Delta X_{0}:=[\Delta X_{0,1}\cdots\Delta X_{0,j}\cdots\Delta X_{0,Q}]\in{\mathbb{R}}^{M\times N}\,.$ (1.8)
Moreover, we define
where the superscript indicates that this is the column-wise reduction of the matrix $\overline{X_{0}}$ in (3.4), below.
We may then define the orthoprojector $P=P^{2}=P^{T}$ onto the span of the family $\{\overline{x_{0,j}}\}_{j=1}^{Q}$ which we assume to be linearly independent. We let $P^{\perp}:={\bf 1}-P$.
We consider a shallow network with hidden layer
$x\mapsto\sigma(W_{1}x+b_{1})\,,$ (1.10)
determined by a weight matrix $W_{1}\in{\mathbb{R}}^{M\times M}$, bias vector $b_{1}\in{\mathbb{R}}^{M}$, and ReLU activation function $\sigma$, which acts component-wise as a ramp function $(a)_{+}:=\max\{0,a\}$. We note that the property of $\sigma$ acting component-wise singles out the given coordinate system in ${\mathbb{R}}^{M}$ as distinct.
We define the output layer with the map ${\mathbb{R}}^{M}\rightarrow{\mathbb{R}}^{Q}$ given by
$x^{\prime}\mapsto W_{2}x^{\prime}+b_{2}$ (1.11)
where $W_{2}\in{\mathbb{R}}^{Q\times M}$ and $b_{2}\in{\mathbb{R}}^{Q}$. Here, we do not include an activation function.
Then, the ${\mathcal{L}}^{2}$ cost function ${\mathcal{C}}[W_{j},b_{j}]$ is defined as
${\mathcal{C}}[W_{j},b_{j}]:=\sqrt{\frac{1}{N}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}\,$ (1.12)
where $|\cdot|_{{\mathbb{R}}^{Q}}$ denotes the Euclidean norm on ${\mathbb{R}}^{Q}$.
Let $R\in O(M)$ denote an orthogonal matrix that diagonalizes $P$. That is, the orthoprojectors
$P_{R}:=RPR^{T}\;\;\;{\rm and}\;P^{\perp}_{R}:=RP^{\perp}R^{T}$ (1.13)
Letting
$u_{M}:=(1,1,\cdots,1)^{T}\in{\mathbb{R}}^{M}$ (1.14)
Then, we construct an upper bound on the cost function by use of the following weights and biases.
We choose
$W_{1}^{*}=R$ (1.15)
and $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ where
$\sigma(RPX_{0}+P_{R}B_{1}^{*})=RPX_{0}+P_{R}B_{1}^{*}\,.$ (1.17)
$\sigma(RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*})=0\,.$ (1.18)
Passing to the output layer, we require $W_{2}^{*}$ to solve
$W_{2}^{*}R\overline{X_{0}^{red}}=Y\,,$ (1.19)
which yields
$W_{2}^{*}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]PR^{T}\,.$ (1.20)
Here,
${\rm Pen}[\overline{X_{0}^{red}}]:=((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}$ (1.21)
denotes the Penrose inverse of $\overline{X_{0}^{red}}$, which satisfies
${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}\;\;\;,\;\;\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]=P\,.$ (1.22)
Finally, we find that
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,,$ (1.23)
that is, it reverts the translation by $P_{R}b_{1}^{*}$ in the previous layer.
We refer to the shallow network defined with these specific weights and biases $W_{i}^{*},b_{i}^{*}$, $i=1,2$, as the constructively trained network. The parameter
$\delta_{P}:=\sup_{j,i}\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|\,,$ (1.24)
measures the relative size between $\Delta X_{0}$ and $\overline{x_{0,j}}$, $j=1,\dots,Q$; in particular, $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like $\frac{|\Delta x|}{|x|}$, and is scaling invariant under $X_{0}\rightarrow\lambda X_{0}$. We then obtain
$\min_{W_{j},b_{j}}{\mathcal{C}}[W_{j},b_{j}]$$\leq$ ${\mathcal{C}}[W_{i}^{*},b_{i}^{*}]$ (1.25)
$\leq$ $\frac{1}{\sqrt{N}}\|Y{\rm Pen}[\overline{X_{0}^{red}}]P\Delta X_{0}\|_{{\mathcal{L}}^{2}}$
$\leq$ $\|Y\|_{op}\;\delta_{P}\;\,.$
In order to match an arbitrary test input $x\in{\mathbb{R}}^{M}$ with one of the output vectors $y_{j}$, $j\in\{1,\dots,Q\}$, we let for $x\in{\mathbb{R}}^{M}$,
${\mathcal{C}}_{j}[x]:=|W_{2}^{*}\sigma(W_{1}^{*}x+b_{1}^{*})+b_{2}^{*}-y_{j}|$ (1.26)
with the weights and biases of the constructively trained network. Then,
$j^{*}={\rm argmin}_{j}{\mathcal{C}}_{j}[x]$ (1.27)
implies that $x$ matches the $j^{*}$-th output $y_{j^{*}}$.
We define the metric
$d_{\widetilde{W}_{2}}(x,y):=|\widetilde{W}_{2}P(x-y)|_{{\mathbb{R}}^{Q}}\;\;\;{\rm for}\;x,y\in{\rm range}(P)\subset{\mathbb{R}}^{M}$ (1.28)
where $\widetilde{W}_{2}:=W_{2}^{*}R$, on the $Q$-dimensional linear subspace ${\rm range}(P)\subset{\mathbb{R}}^{M}$. In Theorem 3.3, we prove that
${\mathcal{C}}_{j}[x]=d_{\widetilde{W}_{2}}(Px,x_{0,j})\,.$ (1.29)
Therefore, matching an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained network is equivalent to solving the metric minimization problem
on the range of $P$.
In other words, $x\in{\mathbb{R}}^{M}$ is matched with the output $y_{j^{*}}\in{\mathbb{R}}^{Q}$ by determining that $Px$ is closest to $\overline{x_{0,j^{*}}}$ among all $\{\overline{x_{0,j}}\in{\rm range}(P)\subset{\mathbb{R}}^{M}\;|\;j=1,\dots,Q\}$ in the $d_{\widetilde{W}_{2}}$ metric.
In Theorem 3.2, we focus on the special case $M=Q$, and present the explicit construction of a degenerate local minimum of the cost function (more precisely, a weighted version of the cost function), and obtain an improvement on the upper bound (1.25) by a factor $1-C_{0}\delta_{P}^{2}$ for some constant $C_{0}\geq 0$. Here, degeneracy means that for all weights and biases $W_{1},b_{1}$ satisfying (1.17), the corresponding minimum of the cost function attains the same value. In particular, this implies that in this range of weights and biases, (1.25) differs from the sharp value with a relative error of order $O(\delta_{P}^{2})$, when $Q=M$.
In Theorem 3.5, we prove a result for $Q=M$ closely related to Theorem 3.2, but for the case where (1.17) is not satisfied.
We expect our results for the case $Q=M$ to be extensible to $Q<M$, but leave a detailed analysis for future work.
This paper is organized as follows. In Section 2, we give a detailed introduction of the mathematical model describings the shallow network. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 4, Theorem 3.3, which we prove in Section 7, and Theorem 3.5. In Section 6, we provide a detailed description of the constructively trained network.
@@ 2. Definition of the mathematical model
In this section, we give a detailed introduction of the mathematical model in discussion, consisting of a shallow network with input layer, one hidden layer, the output layer, and an ${\mathcal{L}}^{2}$ cost function. Letting $Q\in{\mathbb{N}}$ denote the number of distinct output values of the shallow network, we define the output matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q}$ (2.1)
where $y_{j}\in{\mathbb{R}}^{Q}$ is the $j$-th output vector. We assume that the family $\{y_{j}\}$ is linearly independent, so that $Y$ is invertible.
$x_{0,j,i}\in{\mathbb{R}}^{M}\;\;,\;\;i\in\{1,\dots,N_{j}\}$ (2.2)
denotes the $i$-th training input vector corresponding to the $j$-th output vector $y_{j}$. We define the matrix of all training inputs belonging to $y_{j}$ by
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\,.$ (2.3)
Correspondence of a training input vector to the same output $y_{j}$ defines an equivalence relation, that is, for each $j\in\{1,\dots,Q\}$ we have $x_{0,j,i}\sim x_{0,j,i^{\prime}}$ for any $i,i^{\prime}\in\{1,\dots,N_{j}\}$. Accordingly, $X_{0,j}$ labels the equivalence class of all inputs belonging to the $j$-th output $y_{j}$.
Then, we define the matrix of training inputs
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}$ (2.4)
where $N:=\sum_{j=1}^{Q}N_{j}$.
@@
Moreover, we choose to consider a ReLU activation function
@@
is the standard ramp function.
We define
@@
for the terminal layer.
We also define the extension of the matrix of outputs, $Y^{ext}$, by
@@
with $N_{j}$ copies of the same output column vector $y_{j}$. Clearly, $Y^{ext}$ has full rank $Q$.
Then, we consider the ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function
with
$\|A\|_{{\mathcal{L}}^{2}}\equiv\sqrt{{\rm Tr}(AA^{T})}\,,$ (2.15)
where $A^{T}$ is the transpose of the matrix $A$. ${\mathcal{C}}[W_{j},b_{j}]$ is a function of the weights and biases, while the training inputs in $X_{0}$ are considered to be a fixed set of parameters. Clearly, (2.14) is equivalent to (1.12). We note that we are not introducing an activation function in the terminal layer.
Training this shallow network amounts to finding the optimizers $W_{j}^{**},b_{j}^{**}$ for the minimization problem
Before we present our main results in the next section, we contrast our strategy with the commonly used approach via gradient descent.
2.1. Remarks on gradient descent
To determine the solution to the minimization problem (2.16),
${\mathcal{C}}[W_{j}^{**},b_{j}^{**}]=\min_{W_{j},b_{j}}\frac{1}{\sqrt{N}}\|W_{2}\sigma(W_{1}X_{0}+B_{1})+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,,$ (2.17)
one aims to find the minimizers $W_{j}^{**},b_{j}^{**}$ that parametrize the optimally trained network. Our intention here is to explicitly construct (local) minimizers to the cost.
In contrast, minimization via the gradient descent algorithm is based on the dynamical system $({\mathcal{W}},\Phi_{\tau})$ where
${\mathcal{W}}:={\mathbb{R}}^{M^{2}}\times{\mathbb{R}}^{QM}\times{\mathbb{R}}^{M}\times{\mathbb{R}}^{Q}\;\;\;{\rm with}\;\;\underline{\theta}:=(\vec{W_{1}},\vec{W_{2}},b_{1},b_{2})\in{\mathcal{W}}$ (2.18)
is the space of weights and biases, and $\vec{A}:=(A_{1}^{T},\dots,A_{n}^{T})^{T}\in{\mathbb{R}}^{mn}$ with $A_{j}\in{\mathbb{R}}^{m}$ denoting the $j$-th column vector of the matrix $A\in{\mathbb{R}}^{m\times n}$. $\Phi_{\tau}$ denotes the flow generated by the gradient $G_{{\mathcal{C}}}[\underline{\theta}]:=-\nabla_{\underline{\theta}}{\mathcal{C}}[\underline{\theta}]$ of the cost function (or of its square), that is,
$\partial_{\tau}\underline{\theta}(\tau)=G_{{\mathcal{C}}}[\underline{\theta}(\tau)]\;\;\;,\;\;\underline{\theta}(\tau):=\Phi_{\tau}[\underline{\theta}_{0}]\;\;\;,\;\;\underline{\theta}_{0}\in{\mathcal{W}}\,,$ (2.19)
where $\tau\in{\mathbb{R}}_{+}$. Due to
$\partial_{\tau}{\mathcal{C}}[\underline{\theta}(\tau)]=G_{{\mathcal{C}}}[\underline{\theta}(\tau)]\cdot\partial_{t}\underline{\theta}(\tau)=-|G_{{\mathcal{C}}}[\underline{\theta}(\tau)]|^{2}\leq 0\,,$ (2.20)
the cost function ${\mathcal{C}}[\underline{\theta}(\tau)]$ decreases monotonically along orbits of $\Phi_{\tau}$. This is the basis of the majority of computational approaches to the determination of global minima or sufficiently good local minima. Gradient descent algorithms are highly sensitive to the choice of the initial data $\underline{\theta}_{0}$; the latter is in practice mostly picked at random, and vaguely informed by structural insights into the problem.
We distinguish the following situations for the approach via gradient descent.
  1. (1)
    If $W_{1},b_{1}$ are given suitable initial values so that along the gradient descent recursion, $W_{1}X_{0}+B_{1}$ have non-negative matrix components, then the activation function $\sigma$, simply acts as the identity ${\bf 1}_{M\times M}$. Accordingly, (2.17) can be determined by solving the least squares problem
    $\min_{W_{2},b_{2}}\|W_{2}X_{0}+B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,.$ (2.21)
    This is because we can redefine $W_{2}W_{1}\rightarrow W_{2}$ and $b_{2}\rightarrow b_{2}-W_{2}b_{1}$, as $W_{1}$ and $b_{1}$ are unknown. The solution is given by
    $W_{2}X_{0}=Y^{ext}{\mathcal{P}}\;\;\;,\;b_{2}=0$ (2.22)
    where ${\mathcal{P}}^{T}$ is the projector onto the range of $X_{0}^{T}$ (which has rank $Q$), and ${\mathcal{P}}^{\perp}={\bf 1}_{N\times N}-{\mathcal{P}}$.
  2. (2)
    The opposite extreme case is given when $W_{1}$ and $B_{1}$ are given initial values where all matrix components of $W_{1}X_{0}+B_{1}\in{\mathbb{R}}^{M\times N}$ are $\leq 0$, so that
    $\sigma(W_{1}X_{0}+B_{1})=0\,.$ (2.23)
    Then, the cost function becomes independent of $W_{1},b_{1}$, and its gradient in these variables is zero. One is left with the minimization problem
    $\min_{b_{2}}\|B_{2}-Y^{ext}\|_{{\mathcal{L}}^{2}}\,,$ (2.24)
    which has the solution
    $(b_{2})_{i}=\frac{1}{N}\sum_{j=1}^{N}[Y^{ext}]_{ij}=\sum_{j=1}^{Q}\frac{N_{j}}{N}y_{j}\,.$ (2.25)
    The corresponding trained network is not capable of matching inputs $x\in{\mathbb{R}}^{M}$ to outputs $y_{j}$.
  3. (3)
    An intermediate variant of the previous two cases is obtained when $W_{1}$ and $B_{1}$ attain values where some components $(b_{1})_{i_{\ell}}$, $\ell\in{\mathcal{L}}\subset\{1,\dots,M\}$, render the corresponding rows of $[W_{1}X_{0}+B_{1}]_{i_{\ell}j}\leq 0$ for $j=1,\dots,N$. Then, the cost function is independent of $(b_{1})_{i_{\ell}}$ and $[W_{1}]_{i_{\ell}j}$, $j=1,\dots,N$, $\ell\in{\mathcal{L}}$.
    Because by assumption, ${\rm rank}(Y^{ext})=Q$, we find that if
    ${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))=Q\,,$ (2.26)
    the possibility of finding a solution to (2.17) remains intact, based on the surviving components of $W_{1},b_{1}$ that occupy the non-zero rows. In this paper, we will investigate aspects of this scenario.
  4. (4)
    On the other hand, in the same situation as in case (3), if
    ${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))<Q\,,$ (2.27)
    then although one can minimize the cost function depending on the surviving variables, one will not find the global minimum (2.17).
  5. (5)
    Even if the initial values for $W_{1},b_{1}$ lead to case (1), the gradient descent flow might steer them into one of the scenarios in the cases (2) $\sim$ (4) where a rank loss is induced by $\sigma$. Once this occurs, the gradient flow becomes independent of the components of $W_{1},b_{1}$ with respect to which a rank loss is obtained; those components remain stationary (or plateau) at the values at which they are first eliminated by $\sigma$.
We note that a priori, only the bound
${\rm rank}(\sigma(W_{1}X_{0}+B_{1}))\leq{\rm rank}((W_{1}X_{0}+B_{1})\leq M\,,$ (2.28)
is available. In applications, one often finds $M>Q$ or $M\gg Q$. The scenario described in case (3) yields the maximal reduction of complexity to still ensure compatibility with the minimization problem (2.17).
The above discussion highlights the importance of the question whether $W_{1}$ and $b_{1}$ can be analytically determined in a manner that a maximal rank reduction from $M$ to $Q$ is accomplished, as in (2.26), while maintaining the solvability of (2.17). This is addressed in Theorem 3.1, below, based on the construction of accurate upper bounds on the cost function, without making any recourse to gradient flow algorithms.
3. Statement of Main Results
3.1. Upper bound on minimum of cost function for $M>Q$
@@
denote the average over the equivalence class of all training input vectors corresponding to the $j$-th output, and
@@
and assume that its rank is maximal, ${\rm rank}(\overline{X_{0}^{red}})=Q$. The latter condition is necessary for the averages of training inputs to be able to distinguish between the different output values $y_{j}$, $j=1,\dots,Q$.
Because of ${\rm rank}(\overline{X_{0}^{red}})=Q$, the matrix

### v3

@@ Abstract.
In this paper, we approach the problem of cost (loss) minimization in underparametrized shallow ReLU networks through the explicit construction of upper bounds which appeal to the structure of classification data, without use of gradient descent. A key focus is on elucidating the geometric structure of approximate and precise minimizers. We consider an ${\mathcal{L}}^{2}$ cost function, input space ${\mathbb{R}}^{M}$, output space ${\mathbb{R}}^{Q}$ with $Q\leq M$, and training input sample size that can be arbitrarily large. We prove an upper bound on the minimum of the cost function of order $O(\delta_{P})$ where $\delta_{P}$ measures the signal-to-noise ratio of training data. In the special case $M=Q$, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$. The proof of the upper bound yields a constructively trained network; we show that it metrizes a particular $Q$-dimensional subspace in the input space ${\mathbb{R}}^{M}$. We comment on the characterization of the global minimum of the cost function in the given context.
@@ 1. Introduction
Applications of neural networks have rapidly become indispensable in a vast range of research disciplines, and their technological impact in recent years has been enormous. However, despite those successes, the fundamental conceptual reasons underlying their functioning are currently insufficiently understood, and remain the subject of intense investigation, see e.g., [LBH15, Aro18, GK23]. The common approach to the minimization of the cost function is based on the gradient flow generated by it, and a variety of powerful algorithms have been developed for this purpose that can be highly successful in suitable application scenarios; nevertheless, even if a global or sufficiently good local minimum is obtained, the properties of the minimizing parameters (weights, biases) in general remain quite mysterious, see e.g., [HR18, NRS21].
For linear neural networks and the $\mathcal{L}^{2}$ cost, critical points are well understood. Consider a matrix $X\in\mathbb{R}^{M\times N}$, whose columns are made up from the vectors in a training data set, and a matrix $Y\in\mathbb{R}^{Q\times N}$ whose columns are the corresponding target vectors. It is well known that a solution to
$\min_{W}\|WX-Y\|_{\mathcal{L}^{2}}$ (1.1)
is given by $W=Y\,{\rm Pen}[X]$, where ${\rm Pen}[X]$ is the Penrose inverse or pseudoinverse of $X$.
It is helpful to think in terms of projections. Note that ${\rm Pen}[X]X$ is a projector onto the orthogonal complement of the kernel of $X$, ${\rm Pen}[X]X=\mathcal{P}_{(\ker X)^{\perp}}$. Therefore, another way of stating the solution to the linear regression problem is
$\min_{W}\|WX-Y\|_{\mathcal{L}^{2}}=\left\lVert{Y{\rm Pen}[X]X-Y}\right\rVert_{\mathcal{L}^{2}}=\|Y\mathcal{P}_{(\ker X)^{\perp}}^{\perp}\|_{\mathcal{L}^{2}}.$ (1.2)
One can bound this as follows
$\left\lVert{Y\mathcal{P}_{\ker X}}\right\rVert_{\mathcal{L}^{2}}\leq\left\lVert{Y}\right\rVert_{op}\sqrt{\dim\ker X}=\left\lVert{Y}\right\rVert_{op}\sqrt{N-{\rm rank}X}.$ (1.3)
In the overdetermined case $M\geq N$, when $X$ is full-rank the cost can be taken to zero even with a linear network. We assume here that $N$ can be arbitrarily large, so that a zero-loss global minimum is not guaranteed by (1.3).
One of the goals of introducing non-linearities is to increase the complexity of the model. In this work, we study shallow neural networks
$X^{(2)}=W_{2}X^{(1)}+B_{2}\in\mathbb{R}^{Q\times N},\quad X^{(1)}=\sigma(W_{1}X+B_{1})\in\mathbb{R}^{M\times N},$ (1.4)
where $\sigma$ is the source of non-linearity, called an activation function. The ReLU activation function is a popular choice in applications; it is defined on vectors component-wise as the ramp function $(a)_{+}=\max\{0,a\}$. Because it acts as the identity function on vectors with non-negative components, it is convenient to study subcases where the network becomes linear so that results for linear networks apply, and this has been the subject of a few works [Kaw16, ZL18].
Minimizing the $\mathcal{L}^{2}$ cost with respect to $W_{2}$ yields
$\min_{W_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $=\left\lVert{(Y-B_{2}){\rm Pen}[X^{(1)}]X^{(1)}-(Y-B_{2})}\right\rVert_{\mathcal{L}^{2}}$
$=\left\lVert{(Y-B_{2})\mathcal{P}_{\ker X^{(1)}}}\right\rVert_{\mathcal{L}^{2}}.$ (1.5)
The rank of $\mathcal{P}_{\ker X^{(1)}}$ is influenced by the coordinates picked out by $W_{1}$ and the ReLU activation function. When working with classification, there is a natural choice of coordinate frame: Suppose the set of training data is split into $Q$ classes. Then we may consider the mean of each class, and these vectors form the columns of a matrix $\overline{X_{0}^{red}}\in\mathbb{R}^{M\times Q}$. The deviation of each vector in the training data set from the corresponding target class mean is stored in the matrix $\Delta X_{0}$. In Theorem 3.1, we derive an upper bound on the minimum of the cost function with an explicit construction of suitable weights and biases:
$\min_{W_{1},b_{1},W_{2},b_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $\leq C\,\|Y{\rm Pen}[\overline{X_{0}^{red}}]\Delta X_{0}\|_{\mathcal{L}^{2}}.$ (1.6)
Define the parameter
$\delta_{P}:=\sup_{\Delta x_{0}}\left\lvert{{\rm Pen}[\overline{X_{0}^{red}}]\Delta x_{0}}\right\rvert,$ (1.7)
which measures the relative size between the averages in $\overline{X_{0}^{red}}$ and the deviations in $\Delta X_{0}$. Then (1.6) can be bounded in terms of this signal-to-noise ratio,
$\min_{W_{1},b_{1},W_{2},b_{2}}\left\lVert{W_{2}X^{(1)}+B_{2}-Y}\right\rVert_{\mathcal{L}^{2}}$ $\leq C\,\|Y\|_{op}\,\delta_{P}.$ (1.8)
The proof makes crucial use of the bias terms, where $b_{1}$ is constructed so that the terms in the span of the averages survive $\sigma$, and the terms in the complement disappear. When analyzing the particular case $M=Q$, the same strategy creates parameters which result in a linear neural network. Nevertheless, it is still possible to extract a meaningful bound: in Theorem 3.2, we explicitly determine an exact degenerate local minimum of the cost function, and show that the sharp value differs from the upper bound (1.6) obtained for $Q\leq M$ by a relative error $O(\delta_{P}^{2})$.
In Theorem 3.3, we use the parameters given in Theorem 3.1 to define a metric in (a subspace of) input space, and show that the problem of classifying multiclass data is equivalent to a metric minimization problem.
In Theorem 3.5, we prove a result for $Q=M$ closely related to Theorem 3.2, but circumventing the need for the network to be linear, via the introduction of a truncation map.
Finally, a comment on the assumptions: We study shallow neural networks with dimensions of input, hidden and output layer $(M,M,Q)$, respectively, where $M\geq Q$ and at one point specialize to $M=Q$. While this is restrictive in the context of general neural networks, this can be useful as the study of the last layers of a deep network. Moreover, the current paper is the first in a series of works in which we investigate the geometric structure of cost minimizers in neural networks, including deep networks with more general architectures.
1.1. Related work
Despite the power of deep networks, theory of shallow networks is still actively studied. Examples of works on shallow networks include [MVEZ20, ABK+25].
For the ReLU activation function, [Kaw16] and [ZL18] studied minimizers by simplifying to linear networks, with no bias vectors. [Coo21] makes use of the bias vectors to construct a shallow ReLU network which interpolates any data; however, the resulting network is highly overparametrized.
For linear networks, the classic work [BH89] studied critical points of the square loss for shallow linear networks. This was extended in [Kaw16] to deep linear networks, and further studied in works such as [ZL18, TKB19, AMG24].
Remark 1.1.
The inspiration for our approach is drawn from methods in mathematical physics aimed at determining the ground state energy for complex manybody quantum systems (such as large molecules), see for instance [LS10].
@@ 1.1. Related work
We present the explicit construction of local and global ${\mathcal{L}}^{2}$ cost minimizers in underparametrized deep (that is, multilayer) networks in a separate work [CE23], which invokes results developed in the paper at hand, combined with inspiration drawn from rigorous renormalization group analysis in quantum field theory, as for instance in [BCFS03].
1.2. Outline of paper
This paper is organized as follows. In Section 2, we give a detailed description of the mathematical model and introduce notation. In Section 3, we present the main results, namely Theorem 3.1, which we prove in Section 4, Theorem 3.2, which we prove in Section 5, Theorem 3.3, which we prove in Section 6, and Theorem 3.5.
@@ 2. Definition of the mathematical model
In this section, we give a detailed introduction of the mathematical model in discussion, consisting of a shallow neural network for a classification task with $Q$ classes. We also define relevant notation.
@@
Moreover, we choose to consider the ReLU activation function
@@
is the standard ramp function. Then we define the neural network as follows: for a matrix of inputs $X_{0}\in\mathbb{R}^{M\times N}$,
@@
is the terminal layer.
As the labels for the $Q$ classes, we consider a family $\{y_{j}\}_{j=1}^{Q}$ of linearly independent target vectors, and define a target matrix
$Y:=[y_{1},\dots,y_{Q}]\in{\mathbb{R}}^{Q\times Q},$ (2.8)
which is invertible. We also define the extension of the matrix of labels, $Y^{ext}$, by
@@
with $N_{j}$ copies of the same target column vector $y_{j}$. Clearly, $Y^{ext}$ has full rank $Q$.
$x_{0,j,i}\in{\mathbb{R}}^{M}\;\;,\;\;i\in\{1,\dots,N_{j}\}$ (2.11)
denotes the $i$-th training input vector corresponding to the $j$-th target vector $y_{j}$. We define the matrix of all training inputs belonging to $y_{j}$ by
$X_{0,j}:=[x_{0,j,1}\cdots x_{0,j,i}\cdots x_{0,j,N_{j}}]\,.$ (2.12)
Then, we define the matrix of training inputs
$X_{0}:=[X_{0,1}\cdots X_{0,j}\cdots X_{0,Q}]\in{\mathbb{R}}^{M\times N}$ (2.13)
where $N:=\sum_{j=1}^{Q}N_{j}$.
@@
denote the average over the equivalence class of all training input vectors corresponding to the $j$-th class, and
@@
and assume that its rank is maximal, ${\rm rank}(\overline{X_{0}^{red}})=Q$. The latter condition is necessary for the averages of training inputs to be able to distinguish between the different target values $y_{j}$, $j=1,\dots,Q$. Moreover, in this case the matrix